# Good Things Come in Pairs: Paired Autoencoders for Inverse Problems

### v1

@@ 1.5 Latent Space Representation
Figure 3: The top row displays time snapshots from two shallow water equation simulations, demonstrating strikingly different dynamic behaviors. Despite this, both simulations are governed by the same shallow water equations, differing only in their initial conditions, which are displayed in the bottom row.
@@ 1.6 Variational Autoencoders (VAEs)
For a given sample $x\sim\mathcal{X}$, the ELBO takes the form [104, Sect. 4.1], [92, Sect. 21.2.4]
@@ 2.1 Setup
Recent research highlights the effectiveness of paired autoencoders for solving inverse problems [27, 58, 37, 38, 125, 39, 98]. This approach uses latent spaces to compactly represent both data and models as discussed in Section 1.3., leading to inherent regularization and improved inverse solution quality. By mapping between latent spaces, paired autoencoders facilitate an inversion process. This capability underpins a novel hybrid strategy combining data-driven and model-based methods. The result is efficient, regularized solutions achieved without costly forward computations, the difficulties of non-convex optimization, or the computational overhead of explicit regularization parameter tuning.
@@
The latent space mappings $M$ and $M^{\dagger}$ are trainable and parameterize by $\theta_{M}$ and $\theta_{M^{\dagger}}$ respectively, see Figure 4. For instance, one might choose deep neural networks for intricate mappings or use just linear transformations. Non-trainable mappings are also an option, such as the identity mapping, which effectively makes both autoencoders to share the same latent space, i.e., $\mathcal{Z}_{x}=\mathcal{Z}_{y}$. By reducing the trainable parameters through this simplification, we observe that these setups perform effectively in practice, as illustrated in [26]. However, making the latent spaces identical may not always be beneficial for performance, as empirically observed in other experiments, see Section 4.1. Empirically, it has been shown that linear mappings are sufficient to capture the latent space interactions well. Note if we assume, as above, the linear mapping of the autoencoders, we are able to investigate the optimality of the latent space mappings as shown in [58].
@@
$\min_{\theta}$ $\alpha_{x}\mathbb{E}_{x\sim\pi_{x}}\ \mathcal{J}((d_{x}\circ e_{x})(x),x)$ (21)
@@
[…]
Figure 4: Schematic of a paired autoencoder for inverse problems. The $x$-autencoder gives a representation of the QoI, while the $y$-autoencoder provides a representation of the observation space $\mathcal{Y}$. The mappings $M$ and $M^{\dagger}$ between the latent space $\mathcal{Z}_{x}$ and $\mathcal{Z}_{y}$ are surrogate mappings of the forward and the inverse mappings $F$ and $F^{\dagger}$, respectively. A full surrogate forward mapping and inverse mapping $F\approx(d_{y}\circ M\circ e_{x})(x)$ and $F^{\dagger}\approx(d_{x}\circ M^{\dagger}\circ e_{y})(y)$.
@@ 2.2 Reconstruction Quality
A key distinction is made between in-distribution (ID) and out-of-distribution (OOD) samples. ID- samples closely resemble the training data, typically with $x\sim\pi_{x}$ and corresponding $y=F(x)+\varepsilon$, while OOD samples differ significantly. For data-driven approaches, ID samples are expected to be reconstructed accurately. Conversely, OOD samples may and should exhibit poor reconstruction (accurate reconstruction would imply they belong to the training distribution).
@@
Assuming the paired autoencoders have been successfully trained, yielding the encoders $e_{x},e_{y}$, decoders $d_{x},d_{y}$, along with the forward and inverse latent space mappings $M$ and $M^{\dagger}$, respectively. One may want to establish accuracy bounds of the forward prediction $\|F(\hat{x})-y\|$) and of the predicted quantity of interest $\|\hat{x}-x\|$. It has been demonstrated by [27] that these quantities are bounded by computable expressions.
@@
This bound inherently requires all learned mappings $e_{x},e_{y},d_{x},d_{y},M$, and $M^{\dagger}$. We can see that beyond the Lipschitz constants, the tightness of the bound hinges on the precision of the quantity of interest autoencoder $\xi_{x}$, the data autoencoder in the latent space $\xi_{y}$, and how well the latent mappings $\xi_{M}$ can be inverted. Furthermore, $\delta$ is influenced by the accuracy of the forward surrogate and the noise level in the data.
@@
Assuming the forward model $F$ is linear, i.e., $F\in\mathbb{R}^{m\times n}$ and assuming that each of the autoencoder pieces is also linear according to Section 1.4, denoted by $E_{x},E_{y},D_{x}$ and $D_{y}$, we may obtain some theoretical results on the optimality of the latent space mapping. Here we assume $\varepsilon$ is independent noise with zero mean symmetric positive definite covariance $\mathbb{E}(\varepsilon\varepsilon^{\top})=\Gamma_{\varepsilon}=L_{\varepsilon}L_{\varepsilon}^{\top}$.
@@
Let $x$ be a random variable with finite first moment and symmetric positive definite second moment $\Gamma_{x}=L_{x}L_{x}^{\top}$. Further, let $y=Fx+\varepsilon$ with linear forward map $F\in\mathbb{R}^{m\times n}$ and $\varepsilon$ be independent noise with zero mean symmetric positive definite covariance $\mathbb{E}(\varepsilon\varepsilon^{\top})=\Gamma_{\varepsilon}=L_{\varepsilon}L_{\varepsilon}^{\top}$. Given linear autoencoders defined by $E_{x},E_{y},D_{x}$ and $D_{y}$, then if $E_{x}$ has full row-rank, the optimal linear forward mapping between latent spaces with minimal norm is given by
$\hat{M}=E_{y}F\Gamma_{x}E_{x}^{\top}\left(E_{x}\Gamma_{x}E_{x}^{\top}\right)^{-1}\in\operatorname*{arg\,min}_{M}\ \mathbb{E}\left\|MZ_{x}-E_{y}FX\right\|_{2}^{2}.$ (25)
@@
$\hat{M}^{\dagger}=E_{x}\Gamma_{x}^{\top}F^{\top}E_{y}^{\top}\left(E_{y}\Gamma_{y}E_{y}^{\top}\right)^{-1}\in\operatorname*{arg\,min}_{M^{\dagger}}\ \mathbb{E}\left\|M^{\dagger}Z_{y}-Z_{x}\right\|_{2}^{2}.$ (26)
Notice that with these optimal mappings $\hat{M}$ and $\hat{M}^{\dagger}$, surrogate forward and inversion mappings for $F$ and $F^{\dagger}$ are given by $D_{y}\hat{M}E_{x}$ and $D_{x}\hat{M}^{\dagger}E_{y}$, respectively. The proof of Theorem 2.2 can be found in [58].
@@ 2.3 Latent Space Inference
The optional regularization term $\tfrac{\alpha}{2}\left\|z-z_{0}\right\|^{2}$ with $\alpha>0$ controls how much the optimization can change the initial guess
@@ 3.1 Variational Paired Autoencoders
$e_{x}$ $=(\mu_{x;\theta^{e}}(x),\Sigma_{x;\theta^{e}}(x),)$ (31)
$e_{y}$ $=(\mu_{y;\theta^{e}}(y),\Sigma_{y;\theta^{e}}(y)).$ (32)
@@
$z_{x}\sim\mathcal{N}(\mu_{\theta^{e}_{x}}(x),\Sigma_{\theta^{e}}(x)),$ (33)
$z_{y}\sim\mathcal{N}(\mu_{\theta^{e}_{y}}(y),\Sigma_{\theta^{e}}(y)),$ (34)
@@
$(\hat{\mu}_{y},\hat{\Sigma}_{y})$ $=M(\mu_{x;\theta^{e}}(x),\Sigma_{x;\theta^{e}}(x)),$ (37)
$(\hat{\mu}_{x},\hat{\Sigma}_{x})$ $=M^{\dagger}(\mu_{y;\theta^{e}}(y),\Sigma_{y;\theta^{e}}(y)).$ (38)
@@
A simplified illustration is given in Figure 5. While minimizing the above expectations during training, we sample pairs of data and models $\{x\sim\mathcal{X},y\sim\mathcal{Y}\}$ in a stochastic approximation fashion. We also use stochastic approximation using a single sample for the sampling of $z_{x},z_{y},\hat{z}_{x},\hat{z}_{y}$.
[…]
@@ 3.1.1 Inference
Sufficiently many samples enable basic statistical analyses in terms of, e.g., the mean and standard deviation per pixel. The VPAE enables latent space inversion, just like the regular paired autoencoders in Equation 28. However, the VPAE offers more choices for the latent-space initial guess. First, the direct estimate (Equation 44) from the VPAE can serve as the initial guess. Second, each sample from Equation 45 may serve as the initial guess. Third, we could use the sample mean as the initial guess.
@@ 4.1 Image Restoration
In this simulation study, our goal is to reconstruct a clean MNIST image from corrupted versions. We operate under the assumption that the corrupted pixel locations are unknown, so no mask is available. Starting with the clean MNIST $28\times 28$-images $(x_{i})_{i=1}^{60,000}$, we constructed a corrupted dataset $(y_{i})_{i=1}^{60,000}$ by applying a Bernoulli process at the pixel level. For each pixel in each image, we set its value to zero with a probability of $p=0.5$. The resulting training data consisted of pairs of clean and corrupted images $(x_{i},y_{i})_{i=1}^{60,000}$. We then proceeded to establish two autoencoder models, one trained on the clean data distribution, given by $(d_{x}\circ e_{x})(\,\cdot\,)$, and the other on the corrupted data distribution, given by $(d_{y}\circ e_{y})(\,\cdot\,)$. For each of these neural networks, we establish a four-layer CNN with a latent space dimension of $128$ (see Section A.1). Linear latent space mapping $M$ and forward and inverse mappings $M^{\dagger}$ are used. The paired autoencoder networks are trained simultaneously. Reconstruction results for ID test data not seen by the network are shown in the first row of Figure 6, which demonstrate the model’s strong reconstruction performance on ID examples.
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionOriginalCorruptedReconstructedIn DistributionOut of Distribution
@@
As detailed in [27] and further discussed in Section 2.2, we utilize the metrics outlined in Equation 22 to evaluate the likelihood of an image belonging to the training data distribution. Figure 7 displays histograms of these metrics for the ID and OOD samples. While some histograms reveal a distinct difference between the distributions, others show more subtle variations. However, these histograms represent only a one-dimensional projection and thus are useful as a tool to distinguish distributions as opposed to evaluating model performance. A 2D scatter plot further illustrates this distinction (Figure 8). Notably, plotting the first metric from Figure 7 against the second metric clearly reveals a good separation of the clusters. In conclusion, the model’s performance is suboptimal for OOD examples, but the metrics offer a valuable indicator of the expected reconstruction quality for new, unseen data.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption$\frac{\left\|(d_{y}\circ e_{y})(y)-y\right\|_{2}}{\left\|y\right\|_{2}}$$\frac{\left\|(d_{x}\circ e_{x})(\hat{x})-\hat{x}\right\|_{2}}{\left\|\hat{x}\right\|_{2}}$$\frac{\left\|(d_{y}\circ M\circ e_{x})(\hat{x})-y\right\|_{2}}{\left\|y\right\|_{2}}$$\frac{\left\|(M^{\dagger}\circ e_{y})(y)-e_{x}(\hat{x})\right\|_{2}}{\left\|e_{x}(\hat{x})\right\|_{2}}$$\frac{\left\|(M\circ e_{x})(\hat{x})-e_{y}(y)\right\|_{2}}{\left\|e_{y}(y)\right\|_{2}}$

### v2

@@ 1.5 Latent Space Representation
Figure 3: The top row shows snapshots from two different simulations of the shallow water equations taken at the same time point. These images illustrate the water surface elevation (or height) at that instant, highlighting the markedly different dynamical behaviors that emerge from the two cases. Although both simulations evolve under the same shallow water dynamics, the differences arise solely due to variations in their initial conditions, which are depicted in the bottom row. Note that only the elevation field is shown here; the associated velocity fields are not displayed.
@@ 1.6 Variational Autoencoders (VAEs)
For a given sample $x\sim\mathcal{X}$, the ELBO takes the form [118, Sect. 4.1], [105, Sect. 21.2.4]
@@ 2.1 Setup
Recent research highlights the effectiveness of paired autoencoders for solving inverse problems. To the best of our knowledge, the pairing of two autoencoders was first considered by Boink and Brune in [19] and subsequently investigated in [30, 65, 43, 44, 143, 45, 111]. This approach uses latent spaces to compactly represent both data and models as discussed in Section 1.3, leading to inherent regularization and improved inverse solution quality. By mapping between latent spaces, paired autoencoders facilitate an inversion process. This capability underpins a novel hybrid strategy combining data-driven and model-based methods. The result is efficient, regularized solutions achieved without costly forward computations, the difficulties of non-convex optimization, or the computational overhead of explicit regularization parameter tuning.
@@
The latent space mappings $M$ and $M^{\dagger}$ are trainable and parameterize by $\theta_{M}$ and $\theta_{M^{\dagger}}$ respectively, see Figure 4. For instance, one might choose deep neural networks for intricate mappings or use just linear transformations. Non-trainable mappings are also an option, such as the identity mapping, which effectively makes both autoencoders to share the same latent space, i.e., $\mathcal{Z}_{x}=\mathcal{Z}_{y}$. By reducing the trainable parameters through this simplification, we observe that these setups perform effectively in practice, as illustrated in [29]. However, making the latent spaces identical may not always be beneficial for performance, as empirically observed in other experiments, see Section 4.1. Empirically, it has been shown that linear mappings are sufficient to capture the latent space interactions well. Note if we assume, as above, the linear mapping of the autoencoders, we are able to investigate the optimality of the latent space mappings as shown in [65] and stated in Theorem 2.2.
@@
$\min_{\theta}\$$\alpha_{x}\mathbb{E}_{x\sim\pi_{x}}\ \mathcal{J}((d_{x}\circ e_{x})(x),x)$ (25)
@@
$x\begin{matrix}\mbox{encoder}\\[0.0pt] e_{x}\end{matrix}$$z_{x}$$\begin{matrix}\mbox{decoder}\\[0.0pt] d_{x}\end{matrix}$$\tilde{x}$$y\begin{matrix}\mbox{encoder}\\[0.0pt] e_{y}\end{matrix}$$z_{y}$$\begin{matrix}\mbox{decoder}\\[0.0pt] d_{y}\end{matrix}$$\tilde{y}$$MM^{\dagger}$$FF^{\dagger}$
Figure 4: Schematic of a paired autoencoder for inverse problems. The $x$-autoencoder gives a representation of the QoI, while the $y$-autoencoder provides a representation of the observation space $\mathcal{Y}$. The mappings $M$ and $M^{\dagger}$ between the latent space $\mathcal{Z}_{x}$ and $\mathcal{Z}_{y}$ are surrogate mappings of the forward and the inverse mappings $F$ and $F^{\dagger}$, respectively. A full surrogate forward mapping and inverse mapping $F\approx(d_{y}\circ M\circ e_{x})(x)$ and $F^{\dagger}\approx(d_{x}\circ M^{\dagger}\circ e_{y})(y)$.
@@ 2.2 Reconstruction Quality
We categorize data into (i) in-distribution (ID) data, which is drawn from the same distribution as the training set—typically modeled as $x\sim\pi_{x}$ with $y=F(x)+\varepsilon$ – and is expected to be accurately reconstructed if learned properly; (ii) out-of-distribution (OOD) data, which originates from a substantially different distribution and typically lies outside the model’s generalization capacity, and therefore should exhibit poor reconstruction—accurate reconstruction in such cases may suggest a failure to distinguish distributional shift; a third category is near-OOD data, which was not seen during training but lies near the support of the training distribution and may be modeled through mild extrapolation or adaptation [47]. This distinction is crucial for understanding model behavior and robustness, particularly in settings where reconstruction fidelity serves as a proxy for distributional fit.
@@
Assuming the paired autoencoders have been successfully trained, yielding the encoders $e_{x},e_{y}$, decoders $d_{x},d_{y}$, along with the forward and inverse latent space mappings $M$ and $M^{\dagger}$, respectively. One may want to establish accuracy bounds of the forward prediction $\|F(\hat{x})-y\|$ and of the predicted quantity of interest $\|\hat{x}-x\|$. It has been demonstrated in [19, 30] that these quantities are bounded by computable expressions.
@@
The theorem and its proof are provided in [30]. This bound inherently requires all learned mappings $e_{x},e_{y},d_{x},d_{y},M$, and $M^{\dagger}$. We can see that beyond the Lipschitz constants, the tightness of the bound hinges on the precision of the quantity of interest autoencoder $\xi_{x}$, the data autoencoder in the latent space $\xi_{y}$, and how well the latent mappings $\xi_{M}$ can be inverted. Furthermore, $\delta$ is influenced by the accuracy of the forward surrogate and the noise level in the data.
@@
Assuming the forward model $F$ is linear, i.e., $F\in\mathbb{R}^{m\times n}$ and assuming that each of the autoencoder pieces is also linear according to Section 1.4, denoted by $E_{x},E_{y},D_{x}$ and $D_{y}$, we may obtain some theoretical results on the optimality of the latent space mapping. Here we assume $\varepsilon$ is independent noise with zero mean symmetric positive definite covariance ${\mathbb{E}_{\varepsilon\sim\pi_{\varepsilon}}}(\varepsilon\varepsilon^{\top})=\Gamma_{\varepsilon}=L_{\varepsilon}L_{\varepsilon}^{\top}$.
@@
Let $x$ be a random variable with finite first moment and symmetric positive definite second moment $\Gamma_{x}=L_{x}L_{x}^{\top}$. Further, let $y=Fx+\varepsilon$ with linear forward map $F\in\mathbb{R}^{m\times n}$ and $\varepsilon$ be independent noise with zero mean symmetric positive definite covariance ${\mathbb{E}_{\varepsilon\sim\pi_{\varepsilon}}}(\varepsilon\varepsilon^{\top})=\Gamma_{\varepsilon}=L_{\varepsilon}L_{\varepsilon}^{\top}$. Given linear autoencoders defined by $E_{x},E_{y},D_{x}$ and $D_{y}$, then if $E_{x}$ has full row-rank, the optimal linear forward mapping between latent spaces with minimal norm is given by
$\hat{M}=E_{y}F\Gamma_{x}E_{x}^{\top}\left(E_{x}\Gamma_{x}E_{x}^{\top}\right)^{-1}\in\operatorname*{arg\,min}_{M}\ {\mathbb{E}_{x\sim\pi_{x}}\left\|ME_{x}X-E_{y}FX\right\|_{2}^{2}.}$ (29)
@@
$\hat{M}^{\dagger}=E_{x}\Gamma_{x}^{\top}F^{\top}E_{y}^{\top}\left(E_{y}\Gamma_{y}E_{y}^{\top}\right)^{-1}\in\operatorname*{arg\,min}_{M^{\dagger}}\ {\mathbb{E}_{x,\varepsilon\sim\pi_{x},\pi_{\varepsilon}}\left\|M^{\dagger}E_{y}(FX+\varepsilon)-E_{x}X\right\|_{2}^{2}.}$ (30)
Notice that with these optimal mappings $\hat{M}$ and $\hat{M}^{\dagger}$, surrogate forward and inversion mappings for $F$ and $F^{\dagger}$ are given by $D_{y}\hat{M}E_{x}$ and $D_{x}\hat{M}^{\dagger}E_{y}$, respectively. Note that we did not impose further assumptions on the linear forward operator $F$, which may lead to non-uniqueness of $\hat{M}$ and $\hat{M}^{\dagger}$. The proof of Theorem 2.2 can be found in [65].
@@ 2.3 Latent Space Inference
This functional includes neural networks, and we employ automatic differentiation for computation of its gradients. Our numerical experiments use the gradient-based ADAM optimizer [77]. The optional regularization term $\tfrac{\alpha}{2}\left\|z-z_{0}\right\|^{2}$ with $\alpha>0$ controls how much the optimization can change the initial guess
@@
The LSI (32) thus enhances the data-fitting capabilities of the paired autoencoder framework, although at the computational cost of evaluating the forward operator during inference. This moves the computational cost closer to that of the variational or plug-and-play methods.
@@ 3.1 Variational Paired Autoencoders
$e_{x}$ $=(\mu_{x;\theta^{e}_{x}}(x),\Sigma_{x;\theta^{e}_{x}}(x)),$ (35)
$e_{y}$ $=(\mu_{y;\theta^{e}_{y}}(y),\Sigma_{y;\theta^{e}_{y}}(y)).$ (36)
@@
$z_{x}\sim\mathcal{N}(\mu_{x;\theta^{e}_{x}}(x),\Sigma_{x;\theta^{e}_{x}}(x)),$ (37)
$z_{y}\sim\mathcal{N}(\mu_{y;\theta^{e}_{y}}(y),\Sigma_{y;\theta^{e}_{y}}(y)),$ (38)
@@
$(\hat{\mu}_{y},\hat{\Sigma}_{y})$ $=M(\mu_{x;\theta^{e}_{x}}(x),\Sigma_{x;\theta^{e}_{x}}(x)),$ (41)
$(\hat{\mu}_{x},\hat{\Sigma}_{x})$ $=M^{\dagger}(\mu_{y;\theta^{e}_{y}}(y),\Sigma_{y;\theta^{e}_{y}}(y)).$ (42)
@@
A simplified illustration is given in Figure 5. While minimizing the above expectations during training, we sample pairs of data and models $\{x\sim\mathcal{X},y\sim\mathcal{Y}\}$ in a stochastic approximation fashion. We also use stochastic approximation using a single sample for the sampling of $z_{x},z_{y},\hat{z}_{x},\hat{z}_{y}$. We recognize that extending the OOD analysis to the variational paired autoencoder is a promising direction that enables a more rigorous treatment of uncertainty quantification; however, we defer this more complex probabilistic investigation to future work.
$x\begin{matrix}\mbox{encoder}\\[0.0pt] e_{x}\end{matrix}$$z_{x}$$\begin{matrix}\mbox{decoder}\\[0.0pt] d_{x}\end{matrix}$$\tilde{x}$$y\begin{matrix}\mbox{encoder}\\[0.0pt] e_{y}\end{matrix}$$z_{y}$$\begin{matrix}\mbox{decoder}\\[0.0pt] d_{y}\end{matrix}$$\tilde{y}$$MM^{\dagger}$$FF^{\dagger}$Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionKLKL
@@ 3.2 Inference
Sufficiently many samples enable basic statistical analyses in terms of, e.g., the mean and standard deviation per pixel. The VPAE enables latent space inversion, just like the regular paired autoencoders in Equation 32. However, the VPAE offers more choices for the latent-space initial guess. First, the direct estimate (Equation 48) from the VPAE can serve as the initial guess. Second, each sample from 49 may serve as the initial guess. Third, we could use the sample mean as the initial guess.
@@ 4.1 Image Restoration
In this simulation study, our goal is to reconstruct a clean MNIST image from corrupted versions. We operate under the assumption that the corrupted pixel locations are unknown, so no mask is available. Starting with the clean MNIST $28\times 28$-images $(x_{i})_{i=1}^{60,000}$, we constructed a corrupted dataset $(y_{i})_{i=1}^{60,000}$ by applying a Bernoulli process at the pixel level. For each pixel in each image, we set its value to zero with a probability of $p=0.5$. The resulting training data consisted of pairs of clean and corrupted images $(x_{i},y_{i})_{i=1}^{60,000}$. We then proceeded to establish two autoencoder models, one trained on the clean data distribution, given by $(d_{x}\circ e_{x})(\,\cdot\,)$, and the other on the corrupted data distribution, given by $(d_{y}\circ e_{y})(\,\cdot\,)$. For each of these neural networks, we establish a six-layer CNN encoder with a latent space dimension of $128$. The decoder mirrors this structure with additional layers for upsampling (see Appendix 6.1). Linear latent space mapping $M$ and forward and inverse mappings $M^{\dagger}$ are used. The paired autoencoder networks are trained simultaneously. Reconstruction results for ID test data not seen by the network are shown in the first row of Figure 6, which demonstrate the model’s strong reconstruction performance on ID examples.
@@
As detailed in [30] and further discussed in Section 2.2, we utilize the metrics outlined in Section 26 to evaluate the likelihood of an image belonging to the training data distribution. Figure 7 displays histograms of these metrics for the ID and OOD samples. While some histograms reveal a distinct difference between the distributions, others show more subtle variations. However, these histograms represent only a one-dimensional projection and thus are useful as a tool to distinguish distributions as opposed to evaluating model performance. A 2D scatter plot further illustrates this distinction (Figure 8). Notably, plotting the first metric from Figure 7 against the second metric clearly reveals a good separation of the clusters. In conclusion, the model’s performance is suboptimal for OOD examples, but the metrics offer a valuable indicator of the expected reconstruction quality for new, unseen data.